OpenStudy (anonymous):

y+4x=4 (a)(-1,) (b) (4,) find the missing coordinate to complete ordered- pair solution to the given linear equation

OpenStudy (akshay_budhkar):

ok all you need to do is substitute the x co-ordiante in the equation given above to get the y co-ordinate

OpenStudy (akshay_budhkar):

i will do 1 for you, you do the second 1

OpenStudy (akshay_budhkar):

y+4x=4 y+4(-1)=4 y-4=4 y=4+4 y=8

OpenStudy (akshay_budhkar):

you will do the second one right?

OpenStudy (anonymous):

no i need the 2nd one also

OpenStudy (akshay_budhkar):

lol ok i will do it for you...why dont you try though? ( its better if you want to learn, if u just want the answer i will give you u wont benefit though)

OpenStudy (anonymous):

thank you

OpenStudy (akshay_budhkar):

y+4x=4 y+4(4)=4 y+16=4 y=4-16 y=-12.. there you go

OpenStudy (anonymous):

so the ordered pair is(-1, )?

OpenStudy (akshay_budhkar):

(-1,8) that is what i did.. i solved for the y co-ordinate

OpenStudy (anonymous):

and for (4,

OpenStudy (akshay_budhkar):

lol (4,-12)

OpenStudy (anonymous):

thanks

OpenStudy (akshay_budhkar):

my pleasure